OpenStudy (anonymous):

what is the equation of a line that is perpendicular 2x-3y=5 that pases through the point (2,-1)

OpenStudy (anonymous):

Put the equation of the line into slope form by isolating the y. y=2/3x-5/3 To find the slope of a line perpendicular to this one you find the opposite reciprical or in this case: -3/2 Plug in the points you know and the slope into this formula: (y-y1)=m(x-x1) where m=slope x1=x coordinate of a known point y1= y coordinate of a known point in this case we have (y-(-1)=-3/2(x-2) y+1=-3/2x+3 y=-3/2+2

OpenStudy (anonymous):

get it?